(Warranties and Premiums)

WXZ Pools carries a wide variety of residential quality outdoor swimming pools and

supplies to maintain the pools. WXZ uses two sales promotion techniques—warranties and

premiums—to attract and retain customers.

Swimming pools are sold with a two-year warranty for replacement of parts and labor. The

estimated warranty cost, based on past experience, is 6% of sales.

The premium is offered on the purchases of chemical supplies to maintain the pool.

Customers receive a coupon for each dollar spent on pool chemical supplies. Customers

may exchange 250 coupons and $5 for an assortment of pool toys. WXZ pays $16 for each

pool toy and estimates that 40% of the coupons given to customers will be redeemed.

WXZ’s total sales for 2017 were $7,500,000—$6,400,000 from outdoor swimming pools

and $1,100,000 from pool chemicals. Replacement parts and labor for warranty work

Financial Accounting Theory--Chapter 13-Current Liabilities Spring, 2019

12

totaled $234,000 during 2017. A total of 2,500 pool toys used in the premium program

were purchased during the year, and there were 286,000 coupons redeemed in 2017.

The accrual method is used by WXZ to account for the warranty and premium costs for

financial reporting purposes. The balances in the accounts related to warranties and

premiums on January 1, 2017, were as shown below.

Inventory of Premiums

$ 6,100

Premium Liability

11,800

Warranty Liability

149,000

Instructions WXZ Pools is preparing its financial statements for the year ended December

31, 2017. Determine the amounts that will be shown on the 2017 financial statements for

the following.

(1) Warranty Expense.

(2) Warranty Liability.

(3) Premium Expense

(4) Inventory of Premiums.

(5) Premium Liability.

.
Answer #1

1. Warranty Expense = Estimated warranty cost x Swimming pool sales = 6% x $6400000 = $384000

2. Warranty Liability = Beginning balance + Warranty expense - Cost of warranty work = $149000 + $384000 - $234000 = $299000

3. Premium Expense = Estimated pool toys to be given out x Net cost of pool toy = (1100000/250 x 40%) x ($16 - $5) = 1760 x $11 = $19360

4. Inventory of Premiums = Beginning balance + Purchases - Pool toys given out against coupons redeemed = $6100 + (2500 x $16) - (286000/250 x $16) = $6100 + $40000 - $18304 = $27796

5. Premium Liability = Beginning balance + Premium expense - Redeemed = $11800 + $19360 - $12584 = $18576

Redeemed = (286000/250) x ($16 - $5) = 1144 x $11 = $12584